OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. S= 5r^2t, for t A. t= S/5 - r B. t= 25r/S C. t=S/5r^2 D. t= r^2- 5S

OpenStudy (anonymous):

@mathstudent55 ?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@cormacpayne

OpenStudy (anonymous):

\[S = 5r^{2t}\]

OpenStudy (anonymous):

What do i do? /.\

OpenStudy (anonymous):

That's the problem, right? It's not... \[S = 5r^{2}t\] right?

OpenStudy (anonymous):

Yeah :P

OpenStudy (anonymous):

If the problem is ^ that, then it is \[t = \frac{ S }{ 5r^{2} }\]

OpenStudy (anonymous):

But if it is the first one...

OpenStudy (anonymous):

Its the second one :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Okay that makes sense, so it is just a matter of dividing in order to get t by itself on one side

OpenStudy (anonymous):

Can you help me on another one really quick?

OpenStudy (anonymous):

Sure, no problem.

OpenStudy (anonymous):

2|7-7x|= 2x +4 can you solve it? or help me solve it ?

OpenStudy (anonymous):

Sure, so first let's get the absolute value by itself: \[\left| 7 - 7x \right| = x + 2\]

OpenStudy (anonymous):

And now, when you want to get rid of the absolute value, you have to set up two equations: \[7 - 7x = x + 2\] \[7 - 7x = -(x + 2)\]

OpenStudy (anonymous):

Because you want the right side to equal whatever the absolute value could be, which could be either positive or negative. Does that make sense?

OpenStudy (anonymous):

Yeah? lol so i just solve that ^^

OpenStudy (anonymous):

Yup! Solve both equations for x and you have your answer.

OpenStudy (anonymous):

5/-8? :)

OpenStudy (anonymous):

Almost!

OpenStudy (anonymous):

Watch out for your sign changes.

OpenStudy (anonymous):

5/8???

OpenStudy (anonymous):

That's one of the solutions! Now you just have to solve the other equation and you're done!

OpenStudy (anonymous):

5/6?

OpenStudy (anonymous):

3/2?

OpenStudy (anonymous):

Great! Those are your answers!

OpenStudy (anonymous):

Wait so 5/8 and 3/2?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

okay thanks :) If i need help on another can i message you?

OpenStudy (anonymous):

Sure thing!

OpenStudy (anonymous):

okay :)